OpenStudy (anonymous):

Let f(x) = 2x - 6 Solve f-1(x) when x = 2

OpenStudy (anonymous):

is it f^-1 or (f-1)(x)

OpenStudy (anonymous):

if it is f^-1 (x) f(x) = 2x- 6 so we have to find the inverse equation we get y= 2x-6 interchange x and y and make x the subject of the formula x= 2y-6 (x+6)/2 = y our new equation now to find f^-1(x) just insert x which is 2 (2+6)/2 = y 8/2= f^-1(x) 4= f^-1 (x)